OpenStudy (anonymous):

Write the equation y = x – 4 in standard form (Ax + By = C ). A:-2x + y = -2 B: y = -4 C:x + y = 4 D:-x + y = -4

OpenStudy (raden):

subtracted by x to both sides, becomes y = x – 4 y - x = x - 4 - x y - x = -4

OpenStudy (raden):

that's similar with -x + y = -4

OpenStudy (anonymous):

thank you!!

OpenStudy (raden):

welcome :)